OpenStudy (lopezking1):

Solve for y: 3x–4y=8

OpenStudy (anonymous):

first -4Y=-3X+8 then divide by 4 throughout Y= 3/4 X -2

OpenStudy (lopezking1):

whoa wait what lol so is y=3/4x-2 the y=mx+b thing

OpenStudy (anonymous):

they asked for y....so i guess lmao

OpenStudy (lopezking1):

okay that works lol

OpenStudy (e.cociuba):

1.) Isolate the terms with y to the left hand side. 2.) Subtract 3x from both sides:  3.) Solve for y. 4.) Divide both sides by -4: